# Scattering Gravitons off General Spinning Compact Objects to $\mathcal{O}(G^{2}S^{4})$

### v1

@@ Appendix A Decomposing onto a convenient basis
Appendix B The Compton in a general frame
@@ Appendix C Ranked Fourier integrals
$I_{\alpha}^{\mu}(b^{2})=-i\frac{\partial I_{\alpha}(b^{2})}{\partial b^{\mu}}=-\frac{i}{m\omega}\frac{\Gamma(2-\alpha)}{2^{2\alpha+1}\pi\Gamma(\alpha)}\frac{b^{\mu}}{|b|^{4-2\alpha}}\;,$ (64)
$I_{\alpha}^{\mu\nu}(b^{2})=-i\frac{\partial I^{\mu}_{\alpha}(b^{2})}{\partial b^{\nu}}=-\frac{1}{m\omega}\frac{\Gamma(2-\alpha)}{2^{2\alpha+1}\pi\Gamma(\alpha)}\frac{1}{|b|^{6-2\alpha}}\Big[|b|^{2}\Pi^{\mu\nu}+(4-2\alpha)b^{\mu}b^{\nu}\Big]\;,$ (65)
$I_{\alpha}^{\mu\nu\gamma}(b^{2})=-i\frac{\partial I^{\mu\nu}_{\alpha}(b^{2})}{\partial b^{\gamma}}=\frac{i}{m\omega}\frac{\Gamma(2-\alpha)}{2^{2\alpha+1}\pi\Gamma(\alpha)}\frac{(4-2\alpha)}{|b|^{8-2\alpha}}\Big[|b|^{2}(b^{\mu}\Pi^{\nu\gamma}+b^{\nu}\Pi^{\gamma\mu}+b^{\gamma}\Pi^{\mu\nu})+(6-2\alpha)b^{\mu}b^{\nu}b^{\gamma}\Big]\;,$ (66)
$I_{\alpha}^{\mu\nu\gamma\rho}(b^{2})=-i\frac{\partial I^{\mu\nu\gamma}_{\alpha}(b^{2})}{\partial b^{\rho}}=\frac{1}{m\omega}\frac{\Gamma(2-\alpha)}{2^{2\alpha+1}\pi\Gamma(\alpha)}\frac{(4-2\alpha)}{|b|^{10-2\alpha}}\Big[|b|^{4}(\Pi^{\mu\rho}\Pi^{\nu\gamma}+\Pi^{\nu\rho}\Pi^{\gamma\mu}+\Pi^{\gamma\rho}\Pi^{\mu\nu})+(8-2\alpha)(6-2\alpha)b^{\mu}b^{\nu}b^{\gamma}b^{\rho}$

### v2

@@ Appendix A Decomposing onto a convenient basis
Appendix B The Compton amplitude in a general frame
@@ Appendix C Ranked Fourier integrals
$I_{\alpha}^{\mu}(b^{2})=-i\frac{\partial I_{\alpha}(b^{2})}{\partial b_{\mu}}=-\frac{i}{m\omega}\frac{\Gamma(2-\alpha)}{2^{2\alpha+1}\pi\Gamma(\alpha)}\frac{b^{\mu}}{|b|^{4-2\alpha}}\;,$ (64)
$I_{\alpha}^{\mu\nu}(b^{2})=-i\frac{\partial I^{\mu}_{\alpha}(b^{2})}{\partial b_{\nu}}=-\frac{1}{m\omega}\frac{\Gamma(2-\alpha)}{2^{2\alpha+1}\pi\Gamma(\alpha)}\frac{1}{|b|^{6-2\alpha}}\Big[|b|^{2}\Pi^{\mu\nu}+(4-2\alpha)b^{\mu}b^{\nu}\Big]\;,$ (65)
$I_{\alpha}^{\mu\nu\gamma}(b^{2})=-i\frac{\partial I^{\mu\nu}_{\alpha}(b^{2})}{\partial b_{\gamma}}=\frac{i}{m\omega}\frac{\Gamma(2-\alpha)}{2^{2\alpha+1}\pi\Gamma(\alpha)}\frac{(4-2\alpha)}{|b|^{8-2\alpha}}\Big[|b|^{2}(b^{\mu}\Pi^{\nu\gamma}+b^{\nu}\Pi^{\gamma\mu}+b^{\gamma}\Pi^{\mu\nu})+(6-2\alpha)b^{\mu}b^{\nu}b^{\gamma}\Big]\;,$ (66)
$I_{\alpha}^{\mu\nu\gamma\rho}(b^{2})=-i\frac{\partial I^{\mu\nu\gamma}_{\alpha}(b^{2})}{\partial b_{\rho}}=\frac{1}{m\omega}\frac{\Gamma(2-\alpha)}{2^{2\alpha+1}\pi\Gamma(\alpha)}\frac{(4-2\alpha)}{|b|^{10-2\alpha}}\Big[|b|^{4}(\Pi^{\mu\rho}\Pi^{\nu\gamma}+\Pi^{\nu\rho}\Pi^{\gamma\mu}+\Pi^{\gamma\rho}\Pi^{\mu\nu})+(8-2\alpha)(6-2\alpha)b^{\mu}b^{\nu}b^{\gamma}b^{\rho}$